OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points? (-2, -20), (0, -4), (4, -20) Please show all steps ! Will give medal to best answer.

OpenStudy (anonymous):

@Coolsector @Mr.ClayLordMath @ehuman

OpenStudy (anonymous):

find the slope 1st y2 - y1 ------- = slope x2 - x1

OpenStudy (anonymous):

pick 2 of the ordered pairs, remember that they represent (x,y), plug in the values and find slope

OpenStudy (anonymous):

from the slope and one ordered pair you can construct the equation because you will have the x, the y and the slope (m) equation of a line is y=mx+b

OpenStudy (anonymous):

b is the y intercept, or what y value is at x=0

OpenStudy (anonymous):

The slope is 8 I think, @ehuman? So the equation of the line would be -20=8(-2)-20?

OpenStudy (ikram002p):

@ehuman I dont think thats a right answer since parabola means a quadratic equation not leaner as you find it .

OpenStudy (anonymous):

Do you know how to do it the right way @ikram002p?

OpenStudy (jdoe0001):

\(\bf (-2, -20)\qquad (0, -4)\qquad (4, -20)\\ \quad \\ \textit{a parabola equation is a quadratic equation, thus}\qquad y=ax^2+bx+c \) so let us plug in those 3 values given for "x" and "y" in the \(\bf y=ax^2+bx+c\) form, so we'll end up with a system of equations of 3 variables gimme a sec

OpenStudy (jdoe0001):

\(\begin{array}{llll} (-2, -20)\qquad (0, -4)\qquad (4, -20)\qquad ax^2+bx+c\\ \quad \\ (-2, -20)\implies -20 = a(-2)^2+b(-2)+c\implies &\bf4a-2b+c=-20\\ \quad \\ (0, -4)\implies -4= a(0)^2+b(0)+c\implies &\bf c = -4\\ \quad \\ (4, -20)\implies -20= a(4)^2+b(4)+c\implies &\bf 16a+4b+c=-20 \end{array}\)

OpenStudy (jdoe0001):

so based on the 2nd equation, we know c = -4 so we can use that in the 1st and 3rd one and use them like 4a -2b = -24 16a +4b = -24 and from there, all you'd need to do, is find either by elimination or substitution we have "c", already, so just need the "a" and "b" coefficients once you find them, plug them back in at the form \(\bf ax^2+bx+c\)

OpenStudy (anonymous):

Would a be -5.5?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

c is not -5.5 it is -4 as can be seen from the point (0,-4) y = ax^2 +bx + c when x = 0 y=-4 y(0) = c c = -4

OpenStudy (anonymous):

I know c is -4.. I asked if A was -5.5

OpenStudy (anonymous):

follow jdoe0001 steps to find a and b

OpenStudy (anonymous):

I did follow his steps, I wanted to check the values I got..

OpenStudy (anonymous):

4a -2b = -24 16a +4b = -24 if we multiply the first eq by 2 then add it to the second: 24a = -72

OpenStudy (anonymous):

B=3.14??

OpenStudy (anonymous):

a is not -5.5 look at what i wrote

OpenStudy (anonymous):

a=-3

OpenStudy (anonymous):

good.. now find b using one of the equations

OpenStudy (anonymous):

(i mean plug a=-3 into one of the equations and then you will get b)

OpenStudy (ikram002p):

\[y=2x^2+12x-4\] try this

OpenStudy (anonymous):

b=1.71?

OpenStudy (anonymous):

Whats that equation for @ikram002p ?

OpenStudy (anonymous):

no, b should be 6. show me your steps

OpenStudy (ikram002p):

mmmm sry \[y=2x^2-12x-4\]

OpenStudy (anonymous):

16(-3)+4b+4=-20 -48+4b+4=-20 -48=-28b 1.71=b

OpenStudy (ikram002p):

a=2 b=-12 c=-4

OpenStudy (anonymous):

16(-3)+4b+4=-20 -48+4b+4=-20 -48=-28b <- mistake is here why 28b 1.71=b

OpenStudy (anonymous):

16(-3)+4b+4=-20 -48+4b+4=-20 4b = 24 b = 6

OpenStudy (anonymous):

I subtracted 4b from 4b and -24

OpenStudy (anonymous):

Oh I see now. Now I plug it all in to the y=ax^2+bx+c ?

OpenStudy (anonymous):

@Coolsector

OpenStudy (ikram002p):

\[y=ax^2+bx+c\] \[-4=a(0)+b(0)+c \rightarrow c=-4\] \[-20=a(-2)^2+b(-2)-4\rightarrow 4a+2b=16\] \[-20=a(4)^2+b(4)-4\rightarrow 16a+4b=-16\] a=6 &b =-4 the equation would be \[y=6x^2-4x-4\]

OpenStudy (ikram002p):

a=-6 ..

OpenStudy (anonymous):

im sorry but those 2 eq 4a -2b = -24 16a +4b = -24 was incorrect. the correct ones are: 4a-2b=-16 16a+4b=-16 now multiply the first by 2 and add them: 24a = -48 a=-2 now plug a=-2 into the first for example and get: -8-2b=-16 -2b =-8 b =4 so you have y=-2x^2+4x-4 as you can see it pass through all the points: (-2,-20): y(-2) = -8-8-4 = -20 (0,-4) obvious and (4,-20): y(4) = -32+16-4=-20 again, final answer: y=-2x^2+4x-4

OpenStudy (jdoe0001):

hm so much for my miss there \(\bf \begin{array}{llll} (-2, -20)\qquad (0, -4)\qquad (4, -20)\qquad ax^2+bx+c\\ \quad \\ (-2, -20)\implies -20 = a(-2)^2+b(-2)+c\implies &\bf4a-2b+c=-20\\ \quad \\ (0, -4)\implies -4= a(0)^2+b(0)+c\implies &\bf c = -4\\ \quad \\ (4, -20)\implies -20= a(4)^2+b(4)+c\implies &\bf 16a+4b+c=-20 \end{array}\\ \quad \\ 4a-2b-4=-20\\ \quad \\ 16a+4b-4=-20\\ \quad \\ \textit{should be as corrected by Coolsector}\\ \quad \\ 4a-2b=-16\\ \quad \\ 16a+4b=-16\\ \quad \\\)